# Conformalized Late Fusion Multi-View Learning

### v1

@@ 3.2 MVCP: Score Quantile Envelope
$[\beta_{\mathrm{lo}},\beta_{\mathrm{hi}}]\leftarrow[\alpha/M,\alpha],\widehat{\mathcal{Q}}\leftarrow\emptyset,\mathcal{S_{C}}^{(1)}\cup\mathcal{S_{C}}^{(2)}\leftarrow\{(s_{1}(x_{i},y_{i}),...,s_{K}(x_{i},y_{i}))\}_{i=1,N_{\mathcal{C}_{1}}+1}^{N_{\mathcal{C}_{1}},N_{\mathcal{C}_{2}}}$
3:$\{u_{m}\leftarrow\textsc{UnifHypersphere}(K)\}_{m=1}^{M}$
4:while $|\mathcal{S}^{(1)}_{\mathcal{C}}\bigcap\widehat{\mathcal{Q}}|/N_{\mathcal{C}_{1}}\notin 1-\alpha\pm\epsilon$ do
5:     $\beta\leftarrow\frac{\beta_{\mathrm{lo}}+\beta_{\mathrm{hi}}}{2}$
6:     $\left\{\widetilde{q}_{m}\leftarrow\frac{\lceil(N_{\mathcal{C}_{1}}+1)(1-\beta)\rceil}{N_{\mathcal{C}_{1}}}\text{ quantile of }\{u_{m}^{\top}s_{i}\}_{i=1}^{N_{\mathcal{C}_{1}}}\right\}_{m=1}^{M}$
7:     $\widehat{\mathcal{Q}}\leftarrow\bigcap_{m=1}^{M}H(u_{m},\widetilde{q}_{m})$
8:     if $|\mathcal{S}^{(1)}_{\mathcal{C}}\bigcap\widehat{\mathcal{Q}}|/N_{\mathcal{C}_{1}}>1-\alpha$ then $\beta_{\mathrm{lo}}\leftarrow\beta$
9:     else $\beta_{\mathrm{hi}}\leftarrow\beta$
10:$\widehat{t}\leftarrow\frac{\lceil(N_{\mathcal{C}_{2}}+1)(1-\alpha)\rceil}{N_{\mathcal{C}_{2}}}\text{ quantile of }\{\max_{m=1,...,M}(u_{m}^{\top}s_{i}/\widetilde{q}_{m})\}_{i=1}^{N_{\mathcal{C}_{2}}}$
11:Return $\{(u_{m},\widehat{t}\widetilde{q}_{m})\}_{m=1}^{M}$
3.3 MVCP: Prediction Region Definition
@@
In the classification setting, where $|\mathcal{Y}|\in\mathbb{N}$, explicit construction of $\mathcal{C}(x)$ is straightforward. For a fixed $x$, explicitly constructing $\mathcal{C}(x)$ can be done by iterating through $y\in\mathcal{Y}$ and checking if $s(x,y)\in\widehat{\mathcal{Q}}$ by comparing $s(x,y)$ against each one of the defining hyperplanes after projection.
@@
In the case of regression, however, the prediction region cannot be explicitly constructed in the general case, since $\mathcal{Y}$ contains uncountably many elements. In fact, explicit construction is generally not of interest for downstream regression applications. Instead, prediction regions in regression settings are often simply used in their implicit form. We, therefore, focus on one particular application, namely that of [26] discussed in Section 2.2, and demonstrate the MVCP prediction regions can be leveraged in their framework. Such an extension has natural applications to the settings discussed in this original study. For instance, the authors demonstrated their method in a robust traffic routing setting with $c$ being predicted traffic from a probabilistic weather model $q(C\mid X)$ for weather covariates $X$. A multiview application, therefore, naturally emerges with having separately tailored predictive models, such as another $q_{2}(C\mid X_{2})$ predicting traffic based on historical traffic trends.
@@
The original study demonstrated that solving this problem in a computationally efficient manner is feasible by leveraging Danskin’s Theorem so long as $\max_{\widehat{c}\in\mathcal{C}(x)}f(w,\widehat{c})$ is efficiently computable for any fixed $w$. We focus on demonstrating that this remains the case for MVCP, specifically considering the case where individual view score functions take the form of a generalization of the “GPCP” score from [26]. Let $q_{k}(C\mid X_{k})$ represent a generative predictor for view $k$ and $\{\widehat{c}_{jk}\}_{j=1}^{J_{k}}\sim q_{k}(C\mid X_{k})$ be drawn samples. Note that $J_{k}$ need not be constant across $k$. Then, we define the score to be
$s_{k}(x,c)=\min_{j}\left[g_{k}\left(\widehat{c}_{jk},c\right)\right],$ (3)
@@
where $g_{k}(\widehat{c},c)$ is any function that is convex in $c$ for a fixed $\widehat{c}$. Notably this framework subsumes many standard regression settings, e.g., for a deterministic predictor, one can take $q_{k}(C\mid X_{k})=\delta(f_{k}(X_{k}))$. To now compute $\max_{\widehat{c}\in\mathcal{C}(x)}f(w,\widehat{c})$, we first let $\vec{j}\in\mathcal{J}=\{j_{1},...,j_{K}\}$ be an indexing tuple, where each $j_{k}\in\{1,...,J_{k}\}$. That is, each $\vec{j}$ is a vector that “selects” one sample per view. Notably then, the projection $u_{m}^{\top}s(\widehat{c}_{\vec{j}},c)$ is convex in $c$, since the projection directions are all restricted to $\mathcal{S}_{+}^{K-1}$. As a result, we can compute the maximizer $c^{*}$, since
$\begin{gathered}c_{\vec{j}}^{*}:=\operatorname*{arg\,max}_{c}f(w,c)\quad\textrm{s.t.}\quad u_{m}^{\top}s(\widehat{c}_{\vec{j}},c)\leq\widehat{q}_{m}\forall m\in\{1,...,M\}\end{gathered}$ (4)
@@
remains a standard convex optimization problem. The final maximum can then be found by aggregation, namely $c^{*}=\operatorname*{arg\,max}_{\vec{j}\in\mathcal{J}}f(w,c_{\vec{j}}^{*})$. While $|\mathcal{J}|=\prod_{k=1}^{K}J_{k}$, we note that most cases of late-stage fusion in practice tend to have a limited number of views, most typically $K=2$ or $K=3$. This coupled with the fact that computing over these indices is trivially parallelizable means this approach is still computationally tractable. The full procedure is outlined in Algorithm 2.
Algorithm 2 Predict-Then-Optimize Under MVCP Prediction Regions
@@
1:Inputs: Context $x$, CGMs $\{q_{k}(C\mid X_{k})\}_{k=1}^{K}$, Optimization steps $T$, Sample counts $\{J_{k}\}_{k=1}^{K}$, MVCP quantile $\widehat{\mathcal{Q}}:=\{(u_{m},\widehat{q}_{m})\}_{m=1}^{M}$
2:$w\sim U(\mathcal{W}),\{\{\widehat{c}_{jk}\}_{j=1}^{J_{k}}\sim q_{k}(C\mid X_{k})\}_{k=1}^{K},\mathcal{J}=\prod_{k=1}^{K}[J_{k}]$
3:for $t\in\{1,\ldots T\}$ do
@@
4:     $\left\{c_{\vec{j}}^{*}\leftarrow\operatorname*{arg\,max}_{c}\left\{f(w,c)\mid u_{m}^{\top}s(\widehat{c}_{\vec{j}},c)\leq\widehat{q}_{m}\forall m\in\{1,...,M\}\right\}\right\}_{\vec{j}\in\mathcal{J}}$
5:     $c^{*}\leftarrow\operatorname*{arg\,max}_{c_{\vec{j}}^{*}}f(w,c_{\vec{j}}^{*})$
6:     $w\leftarrow\Pi_{\mathcal{W}}(w-\eta\nabla_{w}f(w,c^{*}))$
7:Return $w$
@@ 4 Experiments
We now study MVCP empirically across several tasks, demonstrating its improvement over alternate approaches in a synthetic classification task (Section 4.1), in predict-then-optimize tasks (Section 4.2), and in a classification task over a non-synthetic dataset of interest, namely Handwritten [50] (Section 4.3). To allow for fair comparison, we note that the predictors and calibration and test sets were fixed across choices of calibration procedure for each experiment. This in turn means that some care had to be taken in partitioning $\mathcal{D_{C}}=\mathcal{D_{C}}^{(1)}\cup\mathcal{D_{C}}^{(2)}$ for MVCP, where an insufficiently large $\mathcal{D_{C}}^{(1)}$ would result in poor estimation of the $\alpha$-quantile envelope and hence require a large adjustment $\widehat{t}$ factor and an insufficiently large $\mathcal{D_{C}}^{(2)}$ in the classical reduced predictive efficiency from conformal prediction. For this reason, we fixed the split to be $20\%$-$80\%$ in experiments. Further details are provided in Appendix E, and code is available at: https://github.com/yashpatel5400/fusioncp
4.1 Synthetic Classification Task
We first consider a synthetic classification task, fitting multinomial logistic regression models $\widehat{f}_{k}:\mathcal{X}_{k}\rightarrow\mathcal{Y}$, where $\mathcal{X}=\mathbb{R}^{120}$ and $\mathcal{Y}=\{1,...,20\}$. We consider $K$-view classification problems, with $K\in\{2,...,6\}$, where the views here partition the input dimension into subspaces of dimension $120/K$, namely $\mathcal{X}_{i}=\mathbb{R}^{1+(d/K)(i-1):(d/K)i}$ for $i\in\{1,...,K\}$. We compare MVCP to four baselines: a fusion-free approach that conformalizes the view with the highest predictive accuracy; an approach using a predefined “Norm” fusion algorithm as described in Section 3.1, namely with $g_{1}(s)=(\sum_{k=1}^{K}s_{k}^{2})^{1/2}$; another similarly using a predefined “Sum” fusion algorithm with $g_{2}(s)=\sum_{k=1}^{K}s_{k}$; and one that conformalizes a fused predictor as also discussed in Section 3.1, with $\mathcal{F}=(\widehat{f}_{1}(x)+\dots+\widehat{f}_{K}(x))/K$. We perform conformalization using the standard classification score function across all approaches, namely $s_{k}(x,y)=\sum_{j=1}^{l}\widehat{f}_{k}(x)_{\pi_{j}(x,k)}$ where $y=\pi_{l}(x,k)$ and $\pi(x,k)$ is the permutation of $\{1,\ldots,20\}$ that sorts $\widehat{f}_{k}\left(x\right)$ from most likely to least likely. The results are presented in Table 1.
Table 1: Coverages across tasks for $\alpha=0.10$ are shown in the left table, where coverage was assessed over a batch of 300 i.i.d. test samples. Average prediction set sizes are shown in the right table, averaged over a batch of 300 i.i.d. test samples with standard deviations in parentheses.
$K$ Best View Pre-Fused Sum Norm MVCP
2 0.89 0.91 0.91 0.91 0.90
3 0.89 0.91 0.91 0.91 0.90
4 0.89 0.91 0.90 0.90 0.91
5 0.88 0.90 0.90 0.90 0.90
6 0.89 0.91 0.90 0.90 0.90
Best View Pre-Fused Sum Norm MVCP
10.84 (0.34) 8.11 (0.56) 8.73 (0.42) 8.72 (0.42) 7.58 (0.42)
12.36 (0.30) 8.49 (0.44) 9.05 (0.32) 8.83 (0.35) 7.68 (0.42)
13.32 (0.30) 8.75 (0.51) 9.04 (0.47) 8.41 (0.50) 8.07 (0.50)
13.90 (0.19) 9.01 (0.59) 8.91 (0.63) 8.45 (0.59) 8.37 (0.58)
14.28 (0.27) 9.26 (0.39) 9.04 (0.68) 8.71 (0.71) 8.71 (0.75)
@@ 4.1 Synthetic Classification Task
We see that all the approaches exhibit the desired coverage guarantees. However, the multivariate score approaches consistently produce significantly smaller prediction regions than the individually conformalized input views, with MVCP consistently producing the smallest sets on average.
4.2 MVCP Predict-Then-Optimize
We now study the use of MVCP across predict-then-optimize tasks, specifically considering the fractional knapsack problem as in the original conformal predict-then-optimize paper [26]:
$w^{*}(x):=\min_{w}\max_{\widehat{c}\in\mathcal{C}(X)}\quad-\widehat{c}^{T}w$ (5)

### v2

@@ 3.1.2 Score Quantile Threshold
2:  $[\beta_{\mathrm{lo}},\beta_{\mathrm{hi}}]\leftarrow[\alpha/M,\alpha],\widehat{\mathcal{Q}}\leftarrow\emptyset$
3:  $\mathcal{S_{C}}^{(1)}\cup\mathcal{S_{C}}^{(2)}\leftarrow\{(s_{k}(x_{i},y_{i}))_{k=1}^{K}\}_{i=1,N_{\mathcal{C}_{1}}+1}^{N_{\mathcal{C}_{1}},N_{\mathcal{C}_{2}}}$
4:  $\{u_{m}\leftarrow\textsc{UnifHypersphere}(K)\}_{m=1}^{M}$
5:  while $|\mathcal{S}^{(1)}_{\mathcal{C}}\bigcap\widehat{\mathcal{Q}}|/N_{\mathcal{C}_{1}}\notin 1-\alpha\pm\epsilon$ do
6:     $\beta\leftarrow\frac{\beta_{\mathrm{lo}}+\beta_{\mathrm{hi}}}{2}$
7:     $\left\{\widetilde{q}_{m}\leftarrow(1-\beta)\text{ emp. quantile: }\{u_{m}^{\top}s_{i}\}_{s_{i}\in S_{C}^{(1)}}\right\}_{m=1}^{M}$
8:     $\widehat{\mathcal{Q}}\leftarrow\bigcap_{m=1}^{M}H(u_{m},\widetilde{q}_{m})$
9:     if $|\mathcal{S}^{(1)}_{\mathcal{C}}\bigcap\widehat{\mathcal{Q}}|/N_{\mathcal{C}_{1}}>1-\alpha$ then $\beta_{\mathrm{lo}}\leftarrow\beta$
10:     else $\beta_{\mathrm{hi}}\leftarrow\beta$
11:  end while
12:  $\widehat{t}\leftarrow(1-\alpha)\text{ emp. quantile: }\{\max_{m\in[M]}\frac{u_{m}^{\top}s_{i}}{\widetilde{q}_{m}}\}_{s_{i}\in S_{C}^{(2)}}$
13:  Return $\{(u_{m},\widehat{t}\widetilde{q}_{m})\}_{m=1}^{M}$
3.2 Predict-Then-Optimize
@@
In the classification setting, where $|\mathcal{Y}|\in\mathbb{N}$, explicit construction of $\mathcal{C}(x)$ is straightforward. For a fixed $x$, explicitly constructing $\mathcal{C}(x)$ can be done by iterating through $y\in\mathcal{Y}$ and checking if $s(x,y)\in\widehat{\mathcal{Q}}$ by comparing $s(x,y)$ against each one of the thresholds $\widehat{q}_{m}$ after projection.
@@
In the case of regression, however, the prediction region cannot be explicitly constructed in the general case, since $\mathcal{Y}$ contains uncountably many elements. In fact, explicit construction is generally not of interest for downstream regression applications. Instead, prediction regions in regression settings are often simply used in their implicit form. We, therefore, focus on one particular application, namely that of [36] discussed in Section 2.3, and demonstrate the CSA prediction regions can be leveraged in their framework. Such an extension has natural applications to the settings discussed in this original study. For instance, the authors demonstrated the utility of their method in a robust traffic routing setting with $c$ being predicted traffic from a probabilistic weather model $q(C\mid X)$ for weather covariates $X$. An ensembling approach naturally emerges with having multiple predictive models, such as a $q_{2}(C\mid X)$ predicting traffic based instead on historical trends.
@@
As in Section 3.1, we note that the below described algorithm is better understood with a visual accompaniment, which we provide in Appendix C. [36] demonstrated that solving the robust problem variant $w^{*}(x):=\min_{w}\max_{\widehat{c}\in\mathcal{C}(x)}f(w,\widehat{c})$ in a computationally efficient manner is feasible by performing gradient-based optimization on $w$, where the gradient $\nabla_{w}\phi(w)$ of $\phi(w):=\max_{\widehat{c}\in\mathcal{C}(x)}f(w,\widehat{c})$ can be computed by leveraging Danskin’s Theorem so long as $\max_{\widehat{c}\in\mathcal{C}(x)}f(w,\widehat{c})$ is efficiently computable for any fixed $w$. We focus on demonstrating that this remains the case for CSA, specifically considering the case where individual view score functions take the form of the “GPCP” score considered therein. In this setup, each constituent predictor is a generative model $q_{k}(C\mid X)$ from which $\{\widehat{c}_{kj}\}_{j=1}^{J_{k}}\sim q_{k}(C\mid X)$ samples are drawn. Note that $J_{k}$ need not be constant across $k$. The GPCP score, used to define the score components, is
$s_{k}(x,c)=\min_{j\in 1,...,J_{k}}\left[d_{k}\left(\widehat{c}_{kj},c\right)\right],$ (3)
@@
where $d_{k}(\widehat{c},c)$ is a norm over the $C$ space. Notably, this framework subsumes many standard regression settings, e.g., for a deterministic predictor, one can take $q_{k}(C\mid X)=\delta(f_{k}(X))$. To compute $\max_{\widehat{c}\in\mathcal{C}(x)}f(w,\widehat{c})$, we first let $\vec{j}\in\mathcal{J}=\{j_{1},...,j_{K}\}$ be an indexing tuple, where each $j_{k}\in\{1,...,J_{k}\}$. That is, each $\vec{j}$ is a vector that “selects” one sample per predictor. Notably then, the projection $u_{m}^{\top}s(\widehat{c}_{\vec{j}},c)$ is convex in $c$, since the projection directions are all restricted to $\mathcal{S}_{+}^{K-1}$. As a result, we can compute the maximizer $c^{*}$, since
$c_{\vec{j}}^{*}$ $:=\operatorname*{arg\,max}_{c}f(w,c)$ (4)
   s.t. $u_{m}^{\top}s(\widehat{c}_{\vec{j}},c)\leq\widehat{q}_{m}\quad\forall m\in\{1,...,M\}$
@@
remains a standard convex optimization problem. The final maximum can then be found by aggregation, namely $c^{*}=\operatorname*{arg\,max}_{\vec{j}\in\mathcal{J}}f(w,c_{\vec{j}}^{*})$. While $|\mathcal{J}|=\prod_{k=1}^{K}J_{k}$, we note that certain cases of ensemble prediction, such as multi-view prediction, tend to have a limited number of predictors in practice, most typically $K=2$ or $K=3$. This coupled with the fact that computing over these indices is trivially parallelizable means this approach is still computationally tractable. The full procedure is outlined in Algorithm 2.
Algorithm 2 Predict-Then-Optimize Under CSA
@@
1:  Inputs: Context $x$, Predictors $\{q_{k}(C\mid X)\}_{k=1}^{K}$, Optimization steps $T$, Sample counts $\{J_{k}\}_{k=1}^{K}$, CSA quantile $\{(u_{m},\widehat{q}_{m})\}_{m=1}^{M}$
2:  $\{\{\widehat{c}_{kj}\}_{j=1}^{J_{k}}\sim q_{k}(C\mid X)\}_{k=1}^{K},\mathcal{J}=\prod_{k=1}^{K}[J_{k}]$
3:  $w^{(0)}\sim U(\mathcal{W})$
4:  for $t\in\{1,\ldots T\}$ do
@@
5:     for $\vec{j}\in\mathcal{J}$ do $c_{\vec{j}}^{*}\leftarrow\operatorname*{arg\,max}_{c}f(w^{(t)},c)$$\mathrm{s.t.}\quad\forall m\in 1,...,M\qquad u_{m}^{\top}s(\widehat{c}_{\vec{j}},c)\leq\widehat{q}_{m}$
6:     $c^{*}\leftarrow\operatorname*{arg\,max}_{c_{\vec{j}}^{*}}f(w^{(t)},c_{\vec{j}}^{*})$
7:     $w^{(t)}\leftarrow\Pi_{\mathcal{W}}(w^{(t-1)}-\eta\nabla_{w}f(w^{(t-1)},c^{*}))$
8:  end for
9:  Return $w^{(T)}$
Table 1: Average coverages across tasks for $\alpha=0.10$ are shown in the top rows and average prediction set sizes in the bottom rows. Both were assessed over a batch of i.i.d. test samples (15% of the validation set from ImageNet). Standard deviations and means were computed across 10 randomized draws of the calibration and test sets.
Dataset Models ResNet VGG DenseNet VFCP $\mathcal{C}^{M}$ $\mathcal{C}^{R}$ $\mathcal{C}^{U}$ CSA
ImageNet Coverage 0.901 (0.005) 0.902 (0.003) 0.902 (0.003) 0.899 (0.004) 0.938 (0.003) 0.909 (0.004) 0.9 (0.004) 0.9 (0.003)
Size 137.004 (1.98) 136.116 (2.206) 120.096 (2.427) 46.063 (1.089) 87.337 (1.604) 82.746 (1.692) 131.856 (2.378) 34.006 (0.924)
@@ 4 Experiments
We now study CSA empirically across several tasks, demonstrating its coverage guarantees with reduced conservatism. We demonstrate improvements in an ImageNet classification task in Section 4.1, in real-data UCI regression tasks as studied in [29] in Section 4.2, and in a downstream predict-then-optimize task in Section 4.3.
@@
To allow for fair comparison, we note that the predictors and calibration and test sets were fixed across choices of calibration procedure for each experiment. This in turn means that some care had to be taken in partitioning $\mathcal{D_{C}}=\mathcal{D_{C}}^{(1)}\cup\mathcal{D_{C}}^{(2)}$ for CSA, where an insufficiently large $\mathcal{D_{C}}^{(1)}$ would result in poor estimation of the $\alpha$-quantile envelope and hence require a large adjustment $\widehat{t}$ factor and an insufficiently large $\mathcal{D_{C}}^{(2)}$ in the classical reduced predictive efficiency from conformal prediction. For this reason, we fixed the split to be $20\%$-$80\%$ in experiments.
We herein compare against the methods presented in Section 2.4, namely the model selection of [27], the aggregation methods of [25], and the single weighted score projection (VFCP) of [40]. We specifically consider the following aggregation methods from [25]: the standard majority-vote $\mathcal{C}^{M}$, partially randomized thresholding $\mathcal{C}^{R}$, and fully randomized thresholding $\mathcal{C}^{U}$ approaches, fully described in Appendix F. Notably, these methods do not lend themselves for use in the predict-then-optimize setting, so we eliminate them from consideration therein. Similarly, VFCP can only be applied in classification settings; we, thus, do not compare to it across the regression tasks. Code will be made public upon acceptance.
4.1 Classification Tasks
We first study the predictive efficiency of the aforementioned methods on the ImageNet classification task [44]. In particular, an ensemble was constructed from three separately trained deep learning architectures, namely ResNet-50, VGG-11, and DenseNet-121. Conformalization on the individual models was performed using the standard classification score function across all approaches, namely $s(x,y)=\sum_{j=1}^{l}\widehat{f}(x)_{\pi_{j}(x)}$ where $y=\pi_{l}(x)$ and $\pi(x)$ is the permutation of $\{1,\ldots,|\mathcal{Y}|\}$ that sorts $\widehat{f}\left(x\right)$ from most to least likely. Calibration was performed using 85% of the ImageNet test set and assessment of the coverage and interval lengths on the remaining 15%, with 10 trials conducted over randomized draws of these calibration and test sets.
@@ 4.1 Classification Tasks
The results are presented in Table 1. We see that all the approaches exhibit the desired coverage guarantees. However, the CSA score approach consistently produce significantly smaller prediction regions than both the individually conformalized input views and alternate aggregation strategies.
4.2 UCI Regression Tasks
We now similarly study the predictive efficiency of CSA across a suite of regression tasks from the UCI repository [45] as studied by [29]. We specifically consider those tasks for which there are at least 1,000 samples; the complete list of tasks are provided in Appendix D. The data for each task were split with 80% for training, 10% for calibration, and 10% for testing coverage and interval lengths.
The problem setup was replicated from [29], in which four prediction methods were ensembled for each prediction task, namely an OLS model, a LASSO linear model, a random forest, and a multi-layer perceptron. A standard absolute value residual function was used as the score across all methods, namely $s(x,y)=|\widehat{f}(x)-y|$. Prediction intervals could be explicitly constructed for the $\mathcal{C}^{M}$, $\mathcal{C}^{R}$, and $\mathcal{C}^{U}$ aggregation methods. For assessing interval length for CSA, however, a discretized grid $\mathcal{G}_{Y}\subset\mathcal{Y}$ of coarseness $\Delta y$ was considered, and an interval length estimate given by $\mathcal{L}(\mathcal{C}(x))\approx\Delta y\cdot|\{y:y\in\mathcal{G}_{Y},s(x,y)\in\widehat{\mathcal{Q}}\}|$. $10$ trials were conducted for each task, randomized over subset selections of the training, calibration, and test sets. We also present an ablation, labeled “CSA (Single-Stage),” to demonstrate the two-stage calibration described in Section 3.1 is necessary to retain coverage; this single-stage approach does not split $\mathcal{S}_{C}$ and instead directly computes $\{\widehat{q}_{m}\}$ on $\mathcal{S}_{C}$ per Section 3.1.1.
A subset of the results are given in Table 2; the full set of results is deferred to Appendix D due to space limitations. As in the results presented in Section 4.1, we see that CSA retains the coverage guarantees typical of conformal prediction yet produces significantly smaller prediction intervals than both the individual models and the alternate aggregation strategies. We additionally see that the “CSA (Single-Stage)” approach fails to retain coverage, demonstrating the necessity of the two-stage calibration.
Table 2: Average coverages across tasks for $\alpha=0.05$ are shown in the top rows and average prediction set lengths in the bottom rows, both were assessed over a batch of i.i.d. test samples (10% of the datasets). Standard deviations and means were computed across 10 randomized draws of the training, calibration, and test sets. Note that, while the single-stage prediction regions are the smallest, they fail to achieve the desired coverage level and are, thus, precluded from comparison.
Dataset Methods Linear Model LASSO Random Forest Neural Net $\mathcal{C}^{M}$ $\mathcal{C}^{R}$ $\mathcal{C}^{U}$ CSA (Single-Stage) CSA
3droad Coverage 0.947 (0.001) 0.947 (0.001) 0.948 (0.001) 0.95 (0.0) 0.945 (0.001) 0.915 (0.0) 0.948 (0.0) 0.926 (0.001) 0.95 (0.0)
Length 71.583 (0.325) 71.576 (0.331) 5.362 (0.006) 56.298 (0.602) 53.166 (0.482) 29.04 (0.245) 51.158 (0.015) 1.175 (0.036) 1.361 (0.031)
parkinsons Coverage 0.948 (0.002) 0.951 (0.001) 0.935 (0.005) 0.942 (0.001) 0.946 (0.002) 0.878 (0.006) 0.937 (0.005) 0.857 (0.019) 0.961 (0.004)
Length 37.058 (0.599) 37.349 (0.34) 0.876 (0.11) 18.912 (0.931) 17.918 (0.814) 8.738 (0.427) 23.945 (0.039) 0.465 (0.196) 0.588 (0.188)
pol Coverage 0.942 (0.007) 0.944 (0.007) 0.951 (0.002) 0.952 (0.001) 0.936 (0.006) 0.9 (0.009) 0.949 (0.003) 0.889 (0.007) 0.95 (0.003)
Length 97.558 (0.808) 97.476 (1.0) 19.437 (0.374) 41.775 (1.306) 38.033 (1.092) 27.596 (0.77) 63.011 (0.894) 1.357 (0.019) 1.654 (0.032)
pumadyn32nm Coverage 0.932 (0.008) 0.933 (0.008) 0.943 (0.004) 0.924 (0.011) 0.917 (0.008) 0.886 (0.013) 0.935 (0.008) 0.863 (0.018) 0.926 (0.006)
Length 3.865 (0.043) 3.827 (0.061) 1.158 (0.005) 1.145 (0.012) 1.239 (0.011) 1.089 (0.013) 2.54 (0.029) 2.355 (0.124) 2.767 (0.076)
slice Coverage 0.951 (0.001) 0.948 (0.001) 0.948 (0.0) 0.953 (0.003) 0.947 (0.0) 0.912 (0.0) 0.951 (0.002) 0.91 (0.004) 0.95 (0.001)
Length 35.242 (0.099) 35.863 (0.057) 2.749 (0.059) 8.596 (0.412) 8.315 (0.382) 5.378 (0.216) 20.401 (0.162) 1.114 (0.025) 1.388 (0.02)
4.3 CSA Predict-Then-Optimize
We now study the use of CSA for a real-world predict-then-optimize traffic routing task, as studied in the conformal predict-then-optimize paper [36]. In this task, a time series of $T$ previously observed precipitation readings are used to predict future precipitations and, in turn, anticipated traffic, as fully described in Appendix E. We consider the traffic routing problem for a fixed source-target pair $(s,t)$ over the graph of Manhattan, where $|\mathcal{V}|=4584$ and $|\mathcal{E}|=9867$. Formally, such a problem seeks
$w^{*}(x):=\min_{w}\max_{\widehat{c}\in\mathcal{C}(x)}\widehat{c}^{T}w$ (5)